AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
DESPEC
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu Apr 3 19:29:07 2025
Author
*
:
Subject
*
:
<p>Starting a new night shift - All good</p> <p><strong>1am full check:</strong><br /> DSSSD bias & leakage current ok - attachment # 1-2<br /> FEE64 temperatures ok - attachment # 3<br /> ADC data item stats - attachment # 4<br /> Merger ok - Attachement # 5<br /> Tape service - attachement # 6</p> <p>2:30am: AIDA02 FEE stopped working. DESPEC main DAQ was alerted and remobe AIDA from timestitcher while I was restarting AIDA DAQ.</p> <p>2:45am: AIDA02 FEE is back and AIDA has been added to the timestitcher again in the main DAQ.</p> <p>The AIDA runs at the time of the crash to look at are R7_348, 349.</p> <p>All spectra were zeroed at this time and checked.</p> <p><strong>4:00 am full check:</strong><br /> DSSSD bias & leakage current ok - attachment # 7-8<br /> FEE64 temperatures ok - attachment # 9<br /> ADC data item stats - attachment # 10<br /> Merger ok - attachement # 11<br /> Tape service - attachement # 12<br /> Spectra - attachement # 13<br /> ucesb - attachement # 14</p> <pre> <strong>6:00am full check:</strong> DSSSD bias & leakage current ok - attachment # 15-16 FEE64 temperatures ok - attachment # 17 ADC data item stats - attachment # 18 Merger ok - attachement # 19 Tape service - attachement # 20 Spectra - attachement # 21 ucesb - attachement # 22 6:30 the nearline/offline analysis team reported possible noise issue in DSSD#2 and asked if the threshold were changed at any point. I can't find any mention of changes in the elog. I've not done a setup of the all DAQ when AIDA02 fee crashes earlier this morning. I just did a setup of the merger. Thus I don't think the threshold have changed when I stopeed and restarted the DAQ. I can see that currently slow comparator threshod are set to 0xa for most of the DSSD2 fees except for AIDA06 which is set to 0xf. I'm not sure this is the thershold we would need to increase. Something to check during the next shift.</pre> <p>In the UCESB window I can see that the number of decay in DSSD2 is double the number of decay in DSSD1 while the number of implant is similar. Looking at other shift it looks that this was alsready the case in prvious shifts.</p> <p> </p> <p>07.06 analysis of data file S181/R7_457 - attachment 23</p> <p> max deadtime 5.8% aida08, all other FEE64s < 1.4%</p> <p>LEC data rate 1.788M, HEC data rate 4.3k</p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown